# Towards automatizing detection and quantification of intestinal metaplasia: a multi-expert comparative study

### v1

@@ Abstract
Current gastric cancer risk systems are prone to errors since they evaluate a visual estimation of intestinal metaplasia percentages to assign a risk. This study presents an automated method to detect and quantify intestinal metaplasia using deep learning models as well as a comparative analysis with visual estimations of three experienced pathologists. Gastric samples were collected from two different cohorts: 149 asymptomatic volunteers from a region with a high prevalence of GCa in Colombia and 56 patients from a third-level hospital. Deep learning models were selected and trained to classify intestinal metaplasia, and predictions were used to estimate the percentage of intestinal metaplasia and assign the risk score. Results were compared with independent blinded assessments performed by three experienced pathologists. The best-performing deep learning architecture classified intestinal metaplasia with F1-Score of $0.80\pm 0.01$ and AUC of $0.91\pm 0.01$. Among pathologists, inter-observer agreement by a Fleiss’s Kappa score ranged from $0.61$ to $0.75$. In comparison, agreement between the pathologists and the best-performing model ranged from $0.37$ to $0.54$. Deep learning models show potential to detect and quantify the percentage of intestinal metaplasia with greater precision and reproducibility than experienced pathologists. Likewise, estimated risk shows high inter-observer variability when visually assigning the intestinal metaplasia percentage.
@@ Introduction
Currently, the screening protocol is guided by the Sydney System, introduced in 1991 [6] and updated in 1994 as the Updated Sydney System (USS) [7]. This system recommends histological evaluation of at least five gastric biopsies obtained from both the antrum and corpus, three from lesser curvature and two from greater curvature. This biased estimator has been popularized in practice because the incisura angularis has been considered an area with higher risk and therefore has been included as an additional antrum biopsy [8], and some studies have observed premalignant lesions in this region [9]. Overall, Helicobacter pylori, mononuclear cells, loss of glands (atrophy), intestinal metaplasia, inflammation or dysplasia can be found in any of the five collected biopsies. Particular attention has been paid to intestinal metaplasia which has been scored in different categories: 0 (absence), 1 (mild), 2 (moderate), and 3 (severe), a combination of the extent and topographic distribution of microscopic changes [10]. This type of lesion is considered a gastric adaptation to chronic infection with Helicobacter pylori [11] and a pivotal event in GCa progression, described as a “point of no return" [12]. The final step to establish a progression risk to GCa is to assign a stage [13]. Protocols developed to assign this risk are basically the Operative Link for Gastritis Assessment (OLGA) system [14] and the Operative Link on Gastric Intestinal Metaplasia (OLGIM) system [15] (specifically designed to stage the extent and distribution of intestinal metaplasia). Both systems classify patients from stage 0 (lowest risk) to stage IV (highest risk).
@@
Recently, some concerns have been raised about the uncertainty and clinical implications of current gastric risk stratification [16, 17] after the original USS version was introduced. Several publications have reported variability among general pathologists when assigning OLGIM stages [16, 18, 19] and they have documented that these systems may underestimate or overestimate cancer risk [18, 20, 9]. Although some studies suggest that OLGIM may stratify GCa risk [15, 21], there is currently no robust quantitative evidence to support this statement [22, 18]. Therefore, a more precise and reproducible method to quantify these premalignant lesions remains an unmet need [16]. In this context, automatic approaches are particularly appealing as they address many of the limitations associated with visual assessment. Deep learning, in particular, has demonstrated high accuracy for detecting and quantifying patterns in medical images [16], with outstanding results reported in lung cancer [23, 24, 25], breast cancer [26, 27, 28], prostate cancer [29, 30, 31], and gastric cancer [32, 33, 34]. Specifically in gastric cancer, deep learning models have shown improved diagnostic accuracy and reproducibility, while significantly reducing both inter- and intra-observer variability [35].
@@
The current OLGIM risk assessment system, which routinely guides clinical decisions [36, 12], assigns a risk score by heuristically combining visual estimations of intestinal metaplasia from five different gastric biopsy sites. In this context, intestinal metaplasia remains a critical premalignant GCa stage that requires accurate and reproducible quantification [16]. Interestingly, some studies have reported cases classified as OLGIM low-risk who were later diagnosed with gastric cancer [37, 22], while others identified as high-risk never developed cancer and, in some cases, reversed the lesion [38, 39]. In summary, these investigations indicate that subjectively estimated GCa risk is prone to errors and expert-visually-dependent [12].
In this study, an automatic deep learning approach is proposed to accurately detect and quantify intestinal metaplasia in H&E-stained images. Specifically, the main contributions of this study are as follows:
  • Accurate automatic quantification of the intestinal metaplasia percentage by applying and adapting a set of state-of-the-art neural networks.
  • Comparison of state-of-the-art deep learning models and a foundational model against individual assessment performed by three experienced pathologists to estimate the intestinal metaplasia percentage in H&E-stained images.
  • By testing with two independent cohorts, this investigation assesses model robustness and generalization.
The rest of this document is organized as follows: section Methods describes the proposed methodology, including datasets, deep learning models, automatic detection and scoring of intestinal metaplasia and evaluation criteria. Section Results presents the experimental results and their comparison with the expert assessments. Section Discussion provides a discussion of the findings, including their interpretation, limitations, and implications. Finally, section Conclusion summarizes conclusion of the study.
Methods
This study aims to evaluate automatic quantification of the intestinal metaplasia percentage using deep learning models and to compare these assessments with manual estimations of three experienced pathologists. Four state-of-the-art deep learning models, including a foundational model pre-trained in histopathology images, were fine-tuned and evaluated to classify image patches as positive or negative for intestinal metaplasia. Subsequently, predictions of the best-performing model were used to estimate the percentage of intestinal metaplasia and to assign a score for antrum and corpus gastric biopsies, as shown in Figure 1. These automatic assessments were then compared with manual estimations, followed by an inter-observer agreement analysis and quantification of the variability among pathologists.
@@ Methods
Figure 1: Proposed workflow of the automatic intestinal metaplasia classification with deep learning models and automatic scoring using the proportion of intestinal metaplasia in both antrum and corpus biopsies.
Gastric dataset and demographics
@@
Gastric samples from two independent cohorts were digitized. The first cohort consists of 149 asymptomatic volunteers who were selected from an independent study provided by the CIEDYN foundation, a partner of the Urkunina5000 project [40]. This study performed an endoscopic and pathological characterization of gastric disease in 5.000 volunteers from 55 small villages in Nariño, Colombia, a region long observed as being GCa prevalent for the past two decades. Asymptomatic volunteers were men and women aged 307030-70 years, who had lived in these areas for at least the last 10 years and had never undergone an endoscopy procedure, nor been diagnosed with GCa, any other type of cancer, or any known gastric pathology. The second cohort included 56 patients from the Hospital Universitario Nacional de Colombia included in a larger study which was approved by the ethics committee (Act No. 007, Apr 29 2022). Patients were men and women who present symptoms suggestive of gastritis or other forms of gastric discomfort. However, none of them had a previous GCa diagnosis or any other type of lesion.
@@
At least five gastric biopsies were obtained according to the USS (Fig. 2) in both cohorts. Specifically, two biopsies were taken from the antrum ($A_{1}$, $A_{2}$) (the lesser and greater curvature, both within 2cm2\>cm to 3cm3\>cm far from the pylorus), two biopsies from the corpus ($C_{1}$, $C_{2}$) (from the lesser curvature, approximately 4cm4\>cm proximal to the angle and from the middle portion of the greater curvature, about 8cm8\>cm from the cardias) and one biopsy from the incisura angularis ($A_{3}$) (antral-corpus transition zone). All biopsies were fixed in formalin and embedded in paraffin blocks, from which successive sections from the same case were stained with H&E.
@@
Each case was independently assessed by three experienced pathologists following the Operative Link on Gastric Intestinal Metaplasia (OLGIM) system. This system assumes that variations regarding extent and topographical distribution of intestinal metaplasia reflect distinct clinic-pathological scenarios, each corresponding to a specific GCa risk. Pathologists performed blinded evaluations to estimate the proportion of intestinal metaplasia at five standardized biopsy sites: three from the mucosecreting area (two from the antrum and one from the incisura angularis) and two from the oxyntic mucosa (lesser and greater curvature of the corpus). After intestinal metaplasia was detected, a score was assigned as the estimated percentage of glands with goblet cells, ideally assessed in full-thickness (perpendicular) mucosa sections. Each biopsy was independently scored following a four-tier scale: absence of metaplasia = 0% (score 0), mild metaplasia = 1 – 30% (score 1), moderate metaplasia = 31 – 60% (score 2), and severe metaplasia = > 60% (score 3). The scores for the antral biopsies ($A_{1}$, $A_{2}$, $A_{3}$) and the corpus biopsies ($C_{1}$, $C_{2}$) were averaged separately to obtain a composite score for each region. Finally, risk is determined by combining mean scores of the antrum and corpus, as illustrated in Fig. 2.
@@
All cases were manually annotated by the most experienced pathologist, basically delineating those regions affected by intestinal metaplasia. Overall, these regions contain not only glands with intestinal metaplasia but also a non-negligible amount of stroma and surrounding tissues. The pathologist manually annotated intestinal metaplasia regions using a custom web software developed as part of the Program for the Early Detection of Premalignant Lesions and Gastric Cancer in urban, rural and dispersed areas in the Department of Nariño, equipped with basic tools to annotate and visualize WSI. Annotations were saved according to the Web Annotation Data Model standard [41].
@@
The initial step to detect and quantify intestinal metaplasia involves automatic segmentation of gastric glands. In this study, gland segmentation was performed applying a U-Net architecture with a ResNet18 backbone [42], originally trained to segment colorectal glands. It should be strengthen out that colorectal and gastric glands exhibit remarkable morphological similarities [43], particularly by the presence of goblet cells associated with intestinal metaplasia, which by definition corresponds to structural changes resembling those observed in colorectal tissue [44].
The model was pre-trained using a dataset comprising 165 fields of view extracted from 16 H&E-stained whole-slide images from the GlaS database (Gland Segmentation in Colon Histology Images Challenge) [42]. Since gastric and colorectal glands are alike, this architecture processed the 205 WSI from both gastric cohorts. To adapt the model to the gastric domain, the most experienced pathologist manually annotated 2.434 glands from 15 randomly selected cases, including both metaplastic and non-metaplastic glands. This annotated dataset was used to fine-tune the U-Net architecture, achieving a Dice Score of $0.77\pm 0.22$ with a test subset of 730 glands, that is, glands not included in the training phase.
Since the objective of this study is not only to segment glands but also to detect and quantify intestinal metaplasia, the segmented glands were used to define glandular regions. These regions are basically bounding boxes surrounding groups of glands, within which a grid of fields of view is drawn.
@@
Fields of view labeled as metaplasia or control classes were used to train and evaluate a set of state-of-the-art deep learning models. Specifically, four state-of-the-art convolutional neural networks pre-trained with the ImageNet dataset [45], ResNet50 [46], DenseNet121 [47] and ConvNeXtTiny [48], were adapted to classify microscopic gastric fields. ResNet50 demonstrated remarkable performance in intestinal metaplasia classification, with predictions closely matching the assessments of experienced pathologists [49]. DenseNet121 was used as the basis to construct a specialized histopathology framework [50]. ConvNeXtTiny has demonstrated superior performance at identifying anatomical regions of the stomach in endoscopic images [51]. Furthermore, a foundational model, UNI2-h [52], pre-trained with TCGA [53], a large-scale histopathology image dataset, was used as a feature extractor for the classification of intestinal metaplasia.
@@
The previously presented resulting dataset is imbalanced, with an approximate ratio of 9:1, being the metaplasia class the minority. To address this imbalance, an online data augmentation strategy was implemented during training. This strategy involved applying symmetrical rotations of $90^{\circ}$ increments along the horizontal axis to artificially increase sample diversity and mitigate overfitting. Furthermore, class weights were adjusted before training, assigning greater importance to the minority class (metaplasia) in the loss function to compensate for the skewed class distribution.
@@
The 205 cases in both cohorts were independently evaluated by three experienced pathologists following the OLGIM system. Each pathologist estimated the extent of intestinal metaplasia in both antrum and corpus biopsies, assigning a score from 0 to 3. Based on these scores, an OLGIM stage was assigned: stages 0, I or II indicate the lowest risk, while stages III and IV indicate the highest risk. To assess inter-observer agreement, Fleiss’ Kappa coefficient was calculated to measure overall consistency among the three pathologists, and Cohen’s Kappa coefficient was used to assess pairwise agreement. Both metrics range from 0 (no agreement) to 1 (perfect agreement) with higher values indicating greater agreement.
@@ Results
Deep learning models were trained, validated and tested using the first cohort of 149 cases, and externally tested with the second, independent cohort of 56 cases. The classification results, summarized in Table 1, show mean and standard deviation for both internal and external test sets. These metrics were calculated across three experimental runs during the hyperparameter optimization phase and test sets were left aside from the beginning to ensure unbiased evaluation. Under these conditions, ConvNeXtTiny achieved the highest performance when classifying fields of view, with F1-Score 0.80 $\pm$ 0.01, and AUC 0.91 $\pm$ 0.01, for the internal test set, and F1-Score 0.73 $\pm$ 0.01, and AUC 0.83 $\pm$ 0.01 for the external test set. The UNI2-h foundational model, currently considered a state-of-the-art architecture, showed the poorest performance for intestinal metaplasia classification with F1-Score 0.72 $\pm$ 0.01 and AUC 0.84 $\pm$ 0.01, for the internal test set, and F1-Score 0.62 $\pm$ 0.01 and AUC 0.75 $\pm$ 0.01 for the external test set.
@@
The fields of view predicted by the best-performing model were superimposed to the corresponding whole-slide images of both cohorts. Predictions were plotted in a color-code map: blue and red, the ones correctly classified, and orange and purple, the misclassified ones (Fig. 3). These prediction maps highlight in red areas where the model agrees with the expert-annotated regions. Within these annotated regions, areas in purple correspond to those in which the model disagrees with the expert. Interestingly, the model also identifies metaplastic areas (in orange) that have not previously been annotated. In general, most misclassified fields of view were observed within large regions, which often contain a heterogeneous mix of metaplastic glands, control glands, stroma, and foveolar epithelium.
Figure 3: Comparison between annotated and automatically predicted regions. This visualization illustrates differences between regions annotated by the most experienced pathologist (A-D, for internal test samples, and I-L, for external test samples) with the prediction maps of the best performed model (E-H, for internal test samples, and M-P, for external test samples). Model predictions of fields of view are color-coded: correctly predicted as metaplastic are shown in red, and correctly predicted as control in blue. Control regions incorrectly predicted as metaplasia are displayed in orange and metaplastic regions incorrectly predicted as control in purple.
@@
Inter-observer agreement among pathologists and the best-performing deep learning model is presented in Figure 4. The Cohen’s Kappa coefficient shows pairwise inter-observer agreement in both the antrum and corpus, for both internal and external test sets. Given that experts provided an accurate estimation of the intestinal metaplasia percentage, manual-automatic comparison was performed at a finer level by binning the entire range into incremental intervals of 10%, i.e., ten intervals of 10% instead of the usual 0%, 30%, 60% and more. This uniform partition was used to evaluate agreement between manual and automated estimations. Fleiss’ Kappa, which measures overall agreement among all observers, indicated moderate agreement among the three pathologists, with scores of $0.31$ for the antrum and $0.41$ for the corpus. Overall, agreement between the deep learning model and individual pathologists was comparable to that observed between pairs of pathologists. The level of agreement obtained using the homogeneous partition was also consistent with previously reported OLGIM inter-observer agreement in the literature, with Fleiss’ Kappa scores of $0.48$, indicating moderate agreement [54].
Figure 4: Inter-observer agreement among pathologists and the best-performing deep learning model for the internal and external test sets. PT1, PT2 and PT3 stand for the three pathologists, while MODEL refers to the best-performing automated model (ConvNeXtTiny). Subplots in the first column show agreement for the antrum, and subplots in the second column correspond the corpus.
Subplots in Figure 5 show the inter-observer agreement among pathologists, as well as between the pathologists and the best-performing deep learning model, regarding the OLGIM staging. As previously reported in the literature [55], inter-observer agreement tends to increase when the categorical evaluation scale is simplified. In this work, expert agreement improved significantly when using a coarser grading scheme, the OLGIM staging system.
Figure 5: Inter-observer agreement compared with the automatic OLGIM staging on the internal and external test datasets.
@@
Subplots in Figure 6 display box plots illustrating the relationship between the assigned intestinal metaplasia percentage (y-axis) and the corresponding OLGIM stage (x-axis), as determined by consensus among the three pathologists. As observed, expert variability is higher for intermediate stages, when compared with the other stages, probably because these stages are set by a higher number of antrum and corpus combinations, i.e., different OLGIM scores may produce the same stage. As expected, both manual and automatic estimations exhibit an increasing percentage trend of intestinal metaplasia with higher OLGIM stages. However, the automated model shows substantially lower variability across the four stages. As largely described in the literature, pathologists tend to overestimate the extent of the lesion [18], typically assigning higher percentages compared to the automated model.

### v3

# Towards deep-learning based detection and quantification of intestinal metaplasia on digitized gastric biopsies: a multi-expert comparative study

@@ Abstract
Current gastric cancer (GCa) risk systems are prone to errors since they evaluate a visual estimation of intestinal metaplasia percentages in histopathology images of gastric mucosa to assign a risk. This study presents an automated method to detect and quantify intestinal metaplasia using deep convolutional neural networks as well as a comparative analysis with visual estimations of three experienced pathologists. Gastric samples were collected from two different cohorts: 149 asymptomatic volunteers from a region with a high prevalence of GCa in Colombia and 56 patients from a tertiary hospital. Deep learning models were trained to classify intestinal metaplasia, and predictions were used to estimate a percentage of intestinal metaplasia and to assign an adapted OLGIM stage. Atrophy was not assessed because of the limited reproducibility among pathologists. Results were compared with independent blinded metaplastic assessments performed by three graduated pathologists. The best-performing deep learning architecture classified intestinal metaplasia with F1-Score of $0.80\pm 0.01$ and AUC of $0.91\pm 0.01$. Among pathologists, inter-observer agreement by a Fleiss’s Kappa score ranged from $0.20$ to $0.48$. In comparison, agreement between the pathologists and the best-performing model ranged from $0.12$ to $0.35$. Deep learning models show potential to reliably detect and quantify the percentage of intestinal metaplasia, achieving high classification performance. In practice, visual estimation is still the only available method, yet it is marked by considerable inter-observer variability. Deep learning models provide consistent estimates that could help reduce this subjectivity in risk stratification.
@@ Introduction
Currently, the screening protocol is guided by the Sydney System, introduced in 1991 [6] and updated in 1994 as the Updated Sydney System (USS) [7]. This system recommends histological evaluation of at least five gastric biopsies obtained from both the antrum and corpus, three from lesser curvature and two from greater curvature. This estimator has been popularized in practice because the incisura angularis has been considered an area with higher risk and therefore has been included as an additional antrum biopsy [8], and some studies have observed premalignant lesions in this region [9]. Overall, Helicobacter pylori, mononuclear cells, loss of glands (atrophy), intestinal metaplasia, inflammation or dysplasia can be found in any of the five collected biopsies. Particular attention has been paid to intestinal metaplasia which has been scored in different categories: 0 (absence), 1 (mild), 2 (moderate), and 3 (severe), a combination of the extent and topographic distribution of microscopic changes [10]. This type of lesion is considered a gastric adaptation to chronic infection with Helicobacter pylori [11] and a pivotal event in GCa progression, described as a “point of no return" [12]. The final step to establish a progression risk to GCa is to assign a stage [13]. Protocols developed to assign this risk are basically the Operative Link for Gastritis Assessment (OLGA) system [14] and the Operative Link on Gastric Intestinal Metaplasia (OLGIM) system [15] (specifically designed to stage the extent and distribution of intestinal metaplasia). Both systems classify patients from stage 0 (lowest risk) to stage IV (highest risk).
@@
Recently, some concerns have been raised about the uncertainty and clinical implications of current gastric risk stratification [16, 17] after the original USS version was introduced. Several publications have reported variability among general pathologists when assigning OLGIM stages [16, 18, 19] and they have documented that these systems may underestimate or overestimate cancer risk [18, 20, 9]. Although some studies suggest that OLGIM may stratify GCa risk [15, 21], there is currently no robust quantitative evidence to support this statement [22, 18]. Therefore, a more precise and reproducible method to quantify these premalignant lesions remains an unmet need [16]. In this context, automatic approaches are particularly appealing as they address many of the limitations associated with visual assessment. Deep learning, in particular, has demonstrated high accuracy for detecting and quantifying patterns in medical images [16], with outstanding results reported in lung cancer [23, 24, 25], breast cancer [26, 27, 28], prostate cancer [29, 30, 31], and gastric cancer [32, 33, 34]. Specifically in gastric cancer, deep learning models have shown improved diagnostic accuracy and reproducibility in histopathology images, while significantly reducing both inter- and intra-observer variability [35].
@@
The current OLGIM risk assessment system, which routinely guides in treatment of gastric cancers [36, 12], assigns a risk score by heuristically combining visual estimations of intestinal metaplasia from five different gastric biopsy sites. In this context, intestinal metaplasia remains a critical premalignant GCa stage that requires accurate and reproducible quantification [16]. Interestingly, some studies have reported cases classified as OLGIM low-risk who were later diagnosed with gastric cancer [37, 22], while others identified as high-risk never developed cancer and, in some cases, reversed the lesion [38, 39]. In summary, these investigations indicate that subjectively estimated GCa risk is prone to errors and dependent on expert evaluation [12].
In this study, an automatic deep learning approach is proposed to accurately detect and quantify intestinal metaplasia in hematoxylin and eosin (H&E)-stained images. A main contribution consists in applying and adapting several state-of-the-art neural networks to compute the percentage of intestinal metaplasia, and their performance was compared with manual estimations made by three pathologists. Evaluation with two complementary cohorts, one from a high-risk population of Colombian volunteers and another from a tertiary hospital, allowed assessment of classification accuracy and robustness across different clinical contexts.
Materials and Methods
@@
Figure 1: Workflow of the automatic intestinal metaplasia classification with deep learning models and automatic scoring using the proportion of intestinal metaplasia in both antrum and corpus biopsies.
Data Description
@@
Gastric biopsy samples from two independent cohorts were digitized. The first cohort consists of 149 asymptomatic volunteers who were selected from an independent study provided by the CIEDYN foundation, a partner of the Urkunina5000 project [40]. This study performed an endoscopic and pathological characterization of gastric disease in 5.000 volunteers from 55 small villages in Nariño, Colombia, a region long observed as being GCa prevalent for the past two decades. Asymptomatic volunteers were men and women aged 307030-70 years, who had lived in these areas for at least the last 10 years and had never undergone an endoscopy procedure, nor been diagnosed with GCa, any other type of cancer, or any known gastric pathology. The second cohort included 56 patients from the Hospital Universitario Nacional de Colombia included in a larger study which was approved by the ethics committee (Act No. 007, Apr 29 2022). Patients were men and women who present symptoms suggestive of gastritis or other forms of gastric discomfort. However, none of them had a previous GCa diagnosis or any other type of lesion.
Table 1: Demographic characteristics and OLGIM stage distribution for both data cohorts. The external test set corresponds exclusively to the second cohort.
Characteristics All cases (N=205) Training (N=73) Validation (N=32) Internal Test (N=44) External Test (N=56)
Age (years) 53.68 $\pm$ 10.39 52.68 $\pm$ 10.68 56.62 $\pm$ 10.51 53.20 $\pm$ 9.61 64.73 $\pm$ 13.63
Female (%) 79 (53.02%) 44 (60.27%) 13 (40.62%) 22 (50.00%) 38 (67.86%)
OLGIM
      0 74 (36.10%) 16 (21.92%) 7 (21.88%) 10 (22.73%) 41 (73.21%)
      1 74 (36.10%) 31 (42.47%) 14 (43.75%) 21 (47.73%) 8 (14.29%)
      2 34 (16.59%) 14 (19.18%) 7 (21.88%) 8 (18.18%) 5 (8.93%)
      3 14 (6.83%) 9 (12.33%) 2 (6.25%) 3 (6.82%) 0 (0.00%)
      4 9 (4.39%) 3 (4.11%) 2 (6.25%) 2 (4.55%) 2 (3.57%)
@@
Five gastric biopsies were obtained according to the USS (Fig. 2) in both cohorts. Specifically, two biopsies were taken from the antrum ($A_{1}$, $A_{2}$), other two biopsies from the corpus ($C_{1}$, $C_{2}$), and a final biopsy from the incisura angularis ($A_{3}$) (antral-corpus transition zone). All biopsies were fixed in formalin and embedded in paraffin blocks and stained with H&E
@@
Each case was independently assessed by three experienced pathologists following the OLGIM system. This system assumes that variations regarding extent and topographical distribution of intestinal metaplasia reflect distinct clinic-pathological scenarios, each corresponding to a specific GCa risk. Pathologists performed blinded evaluations to estimate the proportion of intestinal metaplasia at five standardized biopsy sites: three from the mucosecreting area (two from the antrum and one from the incisura angularis) and two from the oxyntic mucosa (lesser and greater curvature of the corpus). After intestinal metaplasia was detected, a score was assigned as the estimated percentage of glands with goblet cells, ideally assessed in full-thickness (perpendicular) mucosa sections. Each biopsy was independently scored following a four-tier scale: absence of metaplasia = 0% (score 0), mild metaplasia = 1 – 30% (score 1), moderate metaplasia = 31 – 60% (score 2), and severe metaplasia = > 60% (score 3). The scores for the antral biopsies ($A_{1}$, $A_{2}$, $A_{3}$) and the corpus biopsies ($C_{1}$, $C_{2}$) were averaged separately to obtain a composite score for each region. Finally, risk was determined by combining mean scores of the antrum and corpus, as illustrated in Fig. 2.
@@
All cases were manually annotated by the most experienced pathologist, delineating those regions affected by intestinal metaplasia. Overall, these regions contain not only glands with intestinal metaplasia but also a non-negligible amount of stroma and surrounding tissues. The pathologist manually annotated intestinal metaplasia regions using a custom web software developed as part of the Program for the Early Detection of Premalignant Lesions and Gastric Cancer in urban, rural and dispersed areas in the Department of Nariño, equipped with basic tools to annotate and visualize WSI. Annotations were saved according to the Web Annotation Data Model standard [41].
@@
The initial step to detect and quantify intestinal metaplasia involves automatic segmentation of gastric glands. In this study, gland segmentation was performed applying a U-Net architecture with a ResNet18 backbone [42], originally trained to segment colorectal glands. This choice was motivated by the fact that colorectal and gastric glands exhibit remarkable morphological similarities [43], particularly by the presence of goblet cells associated with intestinal metaplasia, which by definition corresponds to structural changes resembling those observed in colorectal tissue [44].
The model was pre-trained using a dataset comprising 165 fields of view extracted from 16 H&E-stained whole slide images (WSI) from the GlaS database (Gland Segmentation in Colon Histology Images Challenge) [42]. Since gastric and colorectal glands are similar in appearance, this architecture processed the 205 WSI from both gastric cohorts. To adapt the model to the gastric domain, the most experienced pathologist manually annotated 2.434 glands from 15 randomly selected cases, including both metaplastic and non-metaplastic glands. This annotated dataset was used to fine-tune the U-Net architecture, achieving a Dice Score of $0.77\pm 0.22$ with a test subset of 730 glands, that is, glands not included in the training phase.
Since the objective of this study is not only to segment glands but also to detect and quantify intestinal metaplasia, the segmented glands were used to define glandular regions. These regions were defined as bounding boxes surrounding groups of glands, within which a grid of fields of view is drawn.
@@
Fields of view labeled as metaplasia or control classes were used to train and evaluate a set of state-of-the-art deep learning models. Specifically, four state-of-the-art convolutional neural networks pre-trained with the ImageNet dataset [45], ResNet50 [46], DenseNet121 [47] and ConvNeXtTiny [48], were adapted to classify microscopic gastric fields. ResNet50 was selected based on its remarkable performance in intestinal metaplasia classification, with predictions closely matching the assessments of experienced pathologists [49]. DenseNet121 was used as the basis to construct a specialized histopathology framework [50]. ConvNeXtTiny has demonstrated superior performance at identifying anatomical regions of the stomach in endoscopic images [51]. Furthermore, a foundational model, UNI2-h [52], pre-trained with TCGA [53], a large-scale histopathology image dataset, was used as a feature extractor for the classification of intestinal metaplasia.
@@
A significant class imbalance was noted for the dataset, with a ratio 9:1 for non-metaplasia vs metaplasia patches. To address this imbalance, an online data augmentation strategy was implemented during training. This strategy involved applying symmetrical rotations of $90^{\circ}$ increments along the horizontal axis to artificially increase sample diversity and mitigate overfitting. Furthermore, class weights were adjusted before training, assigning greater importance to the minority class (metaplasia) in the loss function to compensate for the skewed class distribution.
@@
Manual estimations were performed by three pathologists following OLGIM guidelines. Pathologists assigned scores approximating the extent and topographic distribution of intestinal metaplasia across the antrum, incisura, and corpus
@@
The 205 cases were independently evaluated by three pathologists following OLGIM guidelines. Pathologists assigned scores approximating the extent and topographic distribution of intestinal metaplasia across the antrum, incisura, and corpus, and assigned a score from 0 to 3. Based on these scores, an OLGIM stage was assigned: stages 0, I or II indicate the lowest risk, while stages III and IV indicate the highest risk. To assess inter-observer agreement, Fleiss’ Kappa coefficient was calculated to measure overall consistency among the three pathologists, and Cohen’s Kappa coefficient was used to assess pairwise agreement. Both metrics range from 0 (no agreement) to 1 (perfect agreement) with higher values indicating greater agreement.
@@ Results
Deep learning models were fine-tuned and evaluated in a binary classification task (presence vs. absence of intestinal metaplasia) using both cohorts. Specifically, models were trained, validated and tested using the first cohort of 149 cases, and externally tested with the second independent cohort of 56 cases. The classification results, summarized in Table 2, show mean and standard deviation for both internal and external test sets. These metrics were calculated across three experimental runs during the hyperparameter optimization phase and test sets were left aside from the beginning to ensure unbiased evaluation. Under these conditions, ConvNeXtTiny achieved the highest performance when classifying fields of view, with F1-Score 0.80 $\pm$ 0.01, and AUC 0.91 $\pm$ 0.01, for the internal test set, and F1-Score 0.73 $\pm$ 0.01, and AUC 0.83 $\pm$ 0.01 for the external test set. The UNI2-h foundational model, showed the poorest performance for intestinal metaplasia classification with F1-Score 0.72 $\pm$ 0.01 and AUC 0.84 $\pm$ 0.01, for the internal test set, and F1-Score 0.62 $\pm$ 0.01 and AUC 0.75 $\pm$ 0.01 for the external test set.
@@
The fields of view predicted by the best-performing model were superimposed to the corresponding whole-slide images of both cohorts. Predictions were plotted in a color-code map: blue and red, the ones correctly classified, and orange and purple, the misclassified ones (Fig. 3). These prediction maps highlight in red areas where the model agrees with the expert-annotated regions. Within these annotated regions, areas in purple correspond to those in which the model disagrees with the expert. Interestingly, the model also identifies metaplastic areas (in orange) that have not previously been annotated. In general, a majority of misclassified fields of view were observed within large regions, which often contain a heterogeneous mix of metaplastic glands, control glands, stroma, and foveolar epithelium.
Figure 3: Representative examples comparing expert annotations and model predictions. For each case, the left subfigures display the whole-slide region with the most experienced pathologist’s annotation (top) and the corresponding model prediction map (bottom). Marked boxes indicate areas of interest, which are shown at higher magnification on the right. The enlarged fields illustrate regions containing metaplastic glands, as well as portions of tissue annotated by the expert where not all structures correspond to intestinal metaplasia. Model predictions are color-coded as follows: correctly predicted metaplasia (red), correctly predicted control (blue), control regions incorrectly predicted as metaplasia (orange), and metaplastic regions incorrectly predicted as control (purple).
@@
Inter-observer agreement among pathologists and the best-performing deep learning model is presented in Figure 4. The Cohen’s Kappa coefficient shows pairwise inter-observer agreement in both the antrum and corpus, for both internal and external test sets. Given that experts provided an accurate estimation of the intestinal metaplasia percentage, manual-automatic comparison was performed at a finer level by binning the entire range into incremental intervals of 10%, i.e., ten intervals of 10% instead of the usual 0%, 30%, 60% and more. This uniform partition was used to evaluate agreement between manual and automated estimations. Inter-observer agreement among pathologists was moderate, with Fleiss’ Kappa = $0.31$ for antrum, and $0.41$ for corpus, therefore, pairwise Cohen’s Kappa ranged 0.200.200.480.48. Agreement between the model and individual pathologists ranged 0.120.120.350.35. Variability of IM percentage estimates was evident, since pathologists tended to assign higher values than the automated model. These results underscore both the visual estimation variability and the relative consistency of the automated approach. In addition, Spearman’s rank-order correlation was computed between the IM percentage estimates of the model and those of each pathologist, showing significant positive associations ($\rho$ = $0.732\pm 0.011$ for OLGIM stage and $\rho$ = $0.893\pm 0.022$ for IM percentage), indicating model prediction are associated with pathologists’ estimation
Figure 4: Inter-observer agreement in the assignment of intestinal metaplasia (IM) percentage intervals (10% bins across the full range) among pathologists and the best-performing deep learning model (ConvNeXtTiny) for the internal (first row) and external (second row) test sets. PT1, PT2 and PT3 correspond to the three pathologists, while MODEL refers to the automated model. Subplots in the first column show agreement for the antrum, and subplots in the second column correspond to the corpus.
Subplots in Figure 5 show the inter-observer agreement among pathologists, as well as between the pathologists and the best-performing deep learning model, regarding the OLGIM staging. Inter-observer agreement can be seen to increase when the categorical evaluation scale is simplified. In this work, expert agreement improved significantly when using a coarser grading scheme, the OLGIM staging system.
Figure 5: Inter-observer agreement in the assignment of OLGIM stages, compared with the automatic staging on the internal and external test datasets.
@@
Expert estimations of IM percentages showed the greatest variability in intermediate OLGIM stages, likely due to the larger number of possible antrum–corpus combinations that can yield the same stage. Both manual and automated assessments demonstrated a consistent increase in IM percentages with higher OLGIM stages. However, the automated model provided substantially lower variability across the four stages, while pathologists tended to assign higher percentages, reflecting a tendency to overestimate lesion extent. These differences are illustrated in Figure 6.